OpenStudy (anonymous):

Please help giving medal to best answer 💕💕💕💕💕❤ the first four terms of a sequence are shown below 8 5 2 -1

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Which of the following functions best defines this sequence? f(1) = 8, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) + 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 3; for n ≥ 1

OpenStudy (dan815):

its decreasing by 3 and it starts at 8

OpenStudy (dan815):

f(1)= will spit out the first number in the list

OpenStudy (dan815):

f(n) will spit out the nth number in the list

OpenStudy (dan815):

what do you think is the answer?

OpenStudy (anonymous):

Ok thank you . so the answer is d ?

OpenStudy (dan815):

yep

OpenStudy (anonymous):

Thanks :)! can you help me with a couple more?

OpenStudy (dan815):

ok ill try

OpenStudy (anonymous):

ok thanks haha For the function y = f(x), what is the ordered pair for the point on the graph when x = p + 1? (x, f(p)) (x, p + 1) (p + 1, f(p + 1)) (p + 1, f(p)) @dan815